vidyo

vidyo link

Action Items:

News:

Scade il mandato di Roberto Tenchini come rappresentante nazionale, successore eletto a Piacenza.

  • Proporre candidature a decano (F.G.) entro 14 ottobre.
  • Indicazioni da Fabrizio: e' un lavoro a tempo pieno che richiede conoscenze dei meccanismi di CMS e INFN. Extremely demanding.
  • Franco's proposal: T.T.
  • Nomination by local representative.

Elezione nuovo spokeperson alla CMS week di Febbraio. Il percorso verso l'elezione e' descritto qui (restricted access) :

https://indico.cern.ch/event/664612/
  • set of nominations by the end of October.
  • Fabrizio's idea: to propose an extension of the duration of the spokeperson mandate (3 years).

Framework e produzione

  • no news

Attivita' di analisi

B physics

  • FR for the p5' on Sept, 20th
  • discussion last week about work that will be done with Roberto on JPsiPhi. Urs sent code and he is cooperative. Possible joint effort with Milano Bicocca. We will push to have EPR counted for this work.

Dark Matter

Dibosoni, VZ risonante (V=W,Z)

hh4b non risonante

  • proceed with bkg bias checks
  • presentation at hbb meeting tomorrow and Padova meeting next Thursday.

Same sign dileptons

  • feedbacks about note after the presentation at EXO meeting
  • they are not convinced by one of the three methods proposed. There is a correlation among variables which is not fully understood.
  • request to repeat the .. analysis to prove that we have more sensitivity.
  • help needed within 1-2 months

Seesaw a 13 TeV

  • CR and FR done. Publication submitted
  • plan for the future: doubling the statistics will not bring to tighter limit, not enough for a new publication. No plan to work in see-saw but they will continue to work on the multi-leptons.

Trigger

  • student worked 3 months at Fermilab with Roberto for AM track trigger.
  • The AM line has been downgraded wrt the other two approaches (FPGA) due to the risk of building a big chip.
  • Study done on a new small ASIC configuration. It remains as back-up solution. Target for Roberto is to close the project asap.

Detector

  • test on the bi-cell are close to the end. four tubes exposed at gif now.
  • no worse performances (was MB1 a peculiar case? additional materials in MB, check if they outgas under irradiation).
  • another board will be inserted in the gif.

EPR

Flavor tagging

A.O.B.:

  • please upload the EPS proceeding on time!
  • IR of tttt submitted. Young people are recommended to read each paper that will come for IR.
  • Oct 30th,31st timing layer workshop
  • this Thursday presentation from student (one slot free). 21st from Nicola P. and HH. 28th no meeting due to CMS week.
  • PHASE2: TDRs for Muon and Barrel upgrade have been submitted tonight. HCAL TDR work started. Yellow report for HL-LHC work will start soon, projects and suggestion are welcome.

-- FrancoSimonetto - 2017-09-11

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2017-09-12 - MartinoDallOsso
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback